# A Survey of LLM-based Automated Program Repair: Taxonomies, Design Paradigms, and Applications

### v1

@@ 5. Agentic Approaches
Agentic frameworks delegate the choice of each repair step to the language model, rather than prescribing a fixed sequence in code. The key distinction is straightforward: when handwritten logic selects the next action, the approach is procedural, whereas a model-driven choice makes it agentic. Within this category, autonomy rises along three tiers. Tool-augmented agents place one model at the center and let LLMs call repository utilities, build or test commands, and even perform graphical actions, while the outer loop stays static (Yang et al., 2024a, b; Wang et al., 2024; Bouzenia et al., 2024; Zhang et al., 2024b; Luo et al., 2025; Liu et al., 2025). LLM-as-judges detach generation from evaluation by adding a critic LLM that scores or ranks patches, guiding an external search without executing code (He et al., 2025; Zhou et al., 2024c; Li et al., 2024b; Yadavally et al., 2025). Self-controlled systems hand one or more LLMs full authority over planning, delegation, and termination, enabling them to tackle function-level or multi-file bugs at the cost of heavier coordination (Tao et al., 2024; Antoniades et al., 2024; Örwall, 2024; Su et al., 2025). The remainder of this section examines these three paradigms in depth.
@@ 5.1. Tool-Augmented Agents
Tool-augmented agents place a frozen LLM at the center of a fixed perceive-think-act loop and expose a curated toolbox to that loop. The skeleton which “observe the state, reason, select a tool, apply it, repeat” is handwritten and never reordered at run time; the LLMs are limited to deciding which provided tool to call or whether to stop. Systems that follow this pattern include SWE-Agent (Yang et al., 2024a), SWE-Agent M (Yang et al., 2024b), OpenHands (Wang et al., 2024), RepairAgent (Bouzenia et al., 2024), AutoCodeRover (Zhang et al., 2024b), LANTERN (Luo et al., 2025), and VulDebugger (Liu et al., 2025), all of which improve over one-shot prompting by letting LLMs choose among grep, compilation, testing, translation, or debugging actions while the outer control flow stays static.
@@
SWE-Agent M extends SWE-Agent by adding a build_and_screenshot tool that compiles the project, launches a headless browser, captures the rendered UI, and feeds the image to GPT-4V for inspection (Yang et al., 2024b). On their built SWE-bench Multimodal benchmark, it increases the fix rate by 2.3 percentage points over the original agent while maintaining the same reasoning loop, indicating that lightweight visual cues can help resolve issues rooted in front-end behavior without redesigning the core architecture.
@@
RepairAgent wraps an LLM inside a finite-state controller that exposes tools for reading files, running tests, compiling code, and editing sources (Bouzenia et al., 2024). The model chooses actions, observes outputs, and loops until tests pass or progress stalls, reaching 164 fixes on the Defects4J v1.2/2.0 benchmark. Tool freedom lets the agent mimic a developer and achieve state-of-the-art accuracy on this suite. Long exploratory sessions consume many tokens, and careful guardrails are required to avoid infinite loops.
LANTERN tackles low-resource languages by translating buggy code into a language where the LLM is stronger, repairing it there, and translating the patch back (Luo et al., 2025). A decision module picks the intermediate language, and iterative testing guides further rounds of translation and refinement. On the xCodeEval (Khan et al., 2024) benchmark, it raised Rust fix rates by over 22% and improved other under-represented languages as well. Cross-language transfer leverages the strengths of LLMs beyond the source language. Accurate round-trip translation is challenging, and the multi-stage process introduces latency and the risk of semantic drift.
@@
In summary, tool-augmented agents demonstrate that granting an LLM selective tool access significantly enhances fix rates compared to procedural scripts. However, every additional reasoning step and external call still follows the same static loop. Future work should consider caching tool outputs, incorporating lightweight heuristics, or adaptively adjusting the reasoning budget to maintain flexibility while reducing the efficiency gap with purely scripted pipelines.
@@ 5.2. LLM-as-Judges
LLMs can serve not only as code generators but also as judges that evaluate candidate patches. This discriminative role, highlighted by He et al. (He et al., 2025), positions them as scalable proxies for human reviewers when assessing whether a change fixes a defect or eliminates a vulnerability. Concrete judge modules now appear in 4 emerging lines of work: differential patch testing for benchmark validation (Wang et al., 2025), automatic patch correctness assessment (Zhou et al., 2024c), security-fix isolation (Li et al., 2024b), and reference-based execution-free critique (Yadavally et al., 2025). Each line integrates an LLM critic into an APR pipeline to filter, rank, or label patches.
LLM4PatchCorrect uses in-context learning with an LLM for code to judge whether a candidate patch is correct by jointly processing bug descriptions, failing tests, execution traces, test coverage, and previously labeled patches (Zhou et al., 2024c). Using StarCoder as the backbone and evaluated on public automatic patch correctness assessment (APCA) datasets (Defects4J (Just et al., 2014) and BEARS (Madeiral et al., 2019)), it achieved 84.4% accuracy and an 86.5% F1 score without requiring any labels from the target repair tool. In this way, LLM4PatchCorrect can reduce the manual labeling effort when building a model to automatically assess the correctness of generated patches of new APR tools. Both the training and test bugs originate from a limited set of projects, so the tool may not generalize well to other software.
VulSifter identifies the lines in a security-related commit that remove a vulnerability while ignoring cosmetic edits (Li et al., 2024b). It feeds the full diff and vulnerability description into an LLM, then applies handcrafted heuristics to discard incidental changes before accepting the patch. After scanning 5.3 million commits from 127 thousand projects, the workflow isolates 11,632 verified security fixes, and models trained on this purified data generalize better than those trained on raw commits. The LLM-rule hybrid yields high precision, but because it makes decisions after the commit, VulSifter cannot issue real-time alerts during an active repair loop.
Execution-Free Critic supplies instant feedback by comparing a candidate patch with a known ground-truth fix without running code or tests (Yadavally et al., 2025). The critic LLM inspects each modified region, reasons about type safety and semantics, and predicts whether the edited program will compile and behave as the reference solution. On a repository-level benchmark, it matched actual test outcomes in 84.8% of cases and achieved a token-level judgment F1 of 0.916, surpassing static metrics and prior heuristics by up to 72%. Its first advantage is speed: removing build and test execution slashes evaluation latency in multi-step repair loops. Second, token-level commentary highlights exactly which parts of the candidate diverge from the oracle, giving search algorithms fine-grained guidance. The method, however, presupposes access to a correct patch, so it mainly fits evaluation or teacher-student training scenarios rather than real-world debugging. Performance also drops when large, cross-file patches overflow the model context, exposing the usual length limits of transformer critics.
PatchDiff (Wang et al., 2025) pioneers call-trace-guided differential patch testing: it identifies functions touched by a candidate patch, prunes surrounding context, and prompts GPT-4o-mini to generate unit tests that pass under the oracle patch but fail under the candidate, thereby acting as an automated referee. Evaluated on the 500-task SWE-bench Verified benchmark and three leading agents (CodeStory (CodeStoryAI contributors, 2025), LearnByInteract (Su et al., 2025), and OpenHands (Wang et al., 2024)), running all developer tests already invalidates 7.8% of “plausible” patches. In contrast, PatchDiff’s tests expose behavioral divergence in another 29.6%, and manual inspection confirms 28.6% of those as incorrect, inflating leaderboard scores by approximately 6.4 percentage points. The study reveals that most false positives originate from weak test suites and subtle “similar-yet-divergent” implementations, indicating that LLM-driven differential testing can rapidly identify semantic errors without requiring the full execution of the regression suite.
@@ 5.3. Self-Controlled System
Self-controlled systems remove the handwritten skeleton altogether. One or more LLMs generate or revise the global plan at run time, spawn new sub-tasks when needed, and decide when to terminate, so the model, not the script, controls the entire outer loop. Representative examples are MAGIS (Tao et al., 2024), which distributes work to four specialized roles; SWE-Search (Antoniades et al., 2024; Örwall, 2024), which couples LLMs with Monte Carlo tree search to debate and apply patches; and Learn-by-Interact (Su et al., 2025), which lets a single model explore an environment, refine its strategy, and distill the trajectory into reusable instructions. With full control, these agents can tackle multi-file or cross-module defects but incur higher coordination costs and safety risks.
@@
SWE-Search combines Monte Carlo Tree Search with a multi-agent control loop in which GPT-4o iteratively searches the repository, plans edits, and debates candidate patches before applying a fix (Antoniades et al., 2024; Örwall, 2024). On SWE-bench Lite, the system raises the solve rate to 31% with GPT-4o and delivers a 23% relative gain across five LLMs, while the Claude 3.5 Sonnet achieves 39% fixes at 2.7 issues per dollar. However, the search depth scales with inference-time computation and has been noted as “probably pretty expensive,” pointing to future work on cost-aware depth control and quicker repository traversal.
@@
Learn-by-Interact is a self-controlled agentic framework that lets an LLM explore a new software or web environment, plan its action sequence, execute those actions, and then retrospectively distill the trajectory into reusable instructions—so the model, not a script, decides every step of the outer loop, satisfying the 5.3 criterion for full autonomy (Su et al., 2025). They fine-tune a Codestral-22B backbone on the synthesized trajectories and also demonstrate a training-free variant in which Claude-3.5 Sonnet retrieves the distilled instructions at run time to guide its decisions. On SWE-bench and other benchmarks, Learn-by-Interact improves in-context learning accuracy by up to 12.2% and supervised fine-tuning performance by 19.5% over best baselines. However, it relies on detailed and accurate environmental documentation; when documentation is sparse or rapidly changing, the data-synthesis phase can yield low-quality trajectories and incur substantial computational overhead.
@@
Self-controlled agents, therefore, deliver the highest flexibility and already outperform simpler agents on public benchmarks. Three lessons emerge. First, explicit role separation or self-generated sub-goals narrows the search space without sacrificing autonomy. Second, stable feedback signals such as full test suites or critical debates prevent blind exploration. Third, careful cost management remains crucial because token budgets, tool latency, and coordination overhead can offset accuracy gains. Continued progress on adaptive stopping rules, incremental context caching, and lightweight self-reflection is essential before these fully autonomous agents can meet the throughput and reliability demands of continuous integration.
5.4. Discussion and Outlook

### v2

@@ 7. Agentic Approaches
Agentic frameworks delegate the choice of each repair step to the language model, rather than prescribing a fixed sequence in code. The key distinction is straightforward: when handwritten logic selects the next action, the approach is procedural, whereas a model-driven choice makes it agentic. Within this category, autonomy rises along three tiers. Tool-augmented agents place one model at the center and let LLMs call repository utilities, build or test commands, and even perform graphical actions, while the outer loop stays static (Yang et al., 2024a, [n. d.]; Wang et al., 2024; Bouzenia et al., 2025; Zhang et al., 2024b; Luo et al., 2025a; Liu et al., 2025). LLM-as-judges detach generation from evaluation by adding a critic LLM that scores or ranks patches, guiding an external search without executing code (He et al., 2025; Zhou et al., 2024c; Li et al., 2024b; Yadavally et al., 2025). Self-controlled systems hand one or more LLMs full authority over planning, delegation, and termination, enabling them to tackle function-level or multi-file bugs at the cost of heavier coordination (Tao et al., 2024; Antoniades et al., [n. d.]; Örwall, 2024; Hongjin et al., 2025). The remainder of this section examines these three paradigms in depth.
@@ 7.1. Tool-Augmented Agents
Tool-augmented agents place a frozen LLM at the center of a fixed perceive-think-act loop and expose a curated toolbox to that loop. The skeleton, which “observe the state, reason, select a tool, apply it, repeat” is handwritten and never reordered at run time; the LLMs are limited to deciding which provided tool to call or whether to stop. Systems that follow this pattern include SWE-Agent (Yang et al., 2024a), SWE-Agent M (Yang et al., [n. d.]), OpenHands (Wang et al., 2024), RepairAgent (Bouzenia et al., 2025), AutoCodeRover (Zhang et al., 2024b), LANTERN (Luo et al., 2025a), and VulDebugger (Liu et al., 2025), all of which improve over one-shot prompting by letting LLMs choose among grep, compilation, testing, translation, or debugging actions while the outer control flow stays static.
@@
SWE-Agent M extends SWE-Agent by adding a build_and_screenshot tool that compiles the project, launches a headless browser, captures the rendered UI, and feeds the image to GPT-4V for inspection (Yang et al., [n. d.]). On their built SWE-bench Multimodal benchmark, it increases the fix rate by 2.3 percentage points over the original agent while maintaining the same reasoning loop, indicating that lightweight visual cues can help resolve issues rooted in front-end behavior without redesigning the core architecture.
@@
RepairAgent wraps an LLM inside a finite-state controller that exposes tools for reading files, running tests, compiling code, and editing sources (Bouzenia et al., 2025). The model chooses actions, observes outputs, and loops until tests pass or progress stalls, reaching 164 fixes on the Defects4J v1.2/2.0 benchmark. Tool freedom lets the agent mimic a developer and achieve state-of-the-art accuracy on this suite. Long exploratory sessions consume many tokens, and careful guardrails are required to avoid infinite loops.
LANTERN tackles low-resource languages by translating buggy code into a language where the LLM is stronger, repairing it there, and translating the patch back (Luo et al., 2025a). A decision module picks the intermediate language, and iterative testing guides further rounds of translation and refinement. On the xCodeEval (Khan et al., 2024) benchmark, it raised Rust fix rates by over 22% and improved other under-represented languages as well. Cross-language transfer leverages the strengths of LLMs beyond the source language. Accurate round-trip translation is challenging, and the multi-stage process introduces latency and the risk of semantic drift.
@@
In summary, tool-augmented agents demonstrate that granting an LLM selective tool access significantly enhances fix rates compared to procedural scripts. However, every additional reasoning step and external call still follow the same static loop. Future work should consider caching tool outputs, incorporating lightweight heuristics, or adaptively adjusting the reasoning budget to maintain flexibility while reducing the efficiency gap with purely scripted pipelines.
@@ 7.2. LLM-as-Judges
LLMs can serve not only as code generators but also as judges that evaluate candidate patches. This discriminative role, highlighted by He et al. (He et al., 2025), positions them as scalable proxies for human reviewers when assessing whether a change fixes a defect or eliminates a vulnerability. In our taxonomy, such judge components fall under the agentic paradigm whenever their scores gate which candidates or search branches are pursued next or when to stop searching. Concrete judge modules now appear in 4 emerging lines of work: differential patch testing for benchmark validation (Wang et al., 2025), tree-search-guided patch evaluation in iterative repair loops (Hu et al., 2025), dual abstention-and-validation funnels (Cambronero et al., 2025), and patch validation (Ruan et al., 2025).
TSAPR instantiates a tree-search APR framework in which an LLM collaborates with Monte Carlo Tree Search to iteratively select, refine, and evaluate candidate patches instead of following a single-path trial-and-error loop (Hu et al., 2025). In this framework, the LLM plays a dual role as generator and judge: it produces new patches with chain-of-thought reasoning, then reuses the same model as an LLM-as-Judge to score failing candidates based on buggy context, test results, execution feedback, and self-reflection traces, while falling back to a Test-as-Judge policy when test suites provide dense coverage. Empirically, TSAPR repairs 201 out of 835 bugs on Defects4J (Just et al., 2014) under a pass@16 budget and 27 out of 79 vulnerabilities on VUL4J (Bui et al., 2022), surpassing prior APR and AVR baselines under substantially smaller patch budgets. These results suggest that combining global search with self-evaluating LLM critics can turn sparse test outcomes into informative rewards, strengthening LLM-as-Judge for complex multi-location and security-sensitive repairs.
Abstain and Validate introduces a dual LLM policy layer on top of an industrial agentic APR system that learns when to attempt a repair and when to surface a candidate patch to developers, aiming to reduce noise and wasted compute at the repository level during program repair (Cambronero et al., 2025). As LLM-as-Judge components, the bug abstention model reads the bug report and predicts the success probability of the underlying agent, deciding to skip low probability bugs. In contrast, the patch validation model generates a textual fix specification and then judges each candidate patch and its trajectory, assigning a correctness score that filters patches even after build and regression tests pass. On 2 additional sets of 198 null pointer exceptions and 50 sanitizer bugs from Google’s codebase, the dual policies raise filtered fail-to-pass@1 on human bugs from 11.29% to 53% and improve filtered accept@1 by up to 24pp on NPE bugs. These results show that wrapping an agentic repair loop with front-end abstention and back-end LLM-as-Judge validation can sharply increase the precision and perceived reliability of APR in production.
SpecRover builds on AutoCodeRover by coupling iterative specification inference with a multi-agent LLM workflow that searches the codebase, synthesizes function-level intent summaries, and uses them to drive patch generation for GitHub issues on SWE-bench (Ruan et al., 2025). In its review stage, a dedicated LLM reviewer agent acts as an LLM-as-Judge, executing the reproducer test on the original and patched programs, inspecting inferred specifications, tests, and issue descriptions, and then either rejecting the patch with natural-language feedback or approving it and triggering regression testing. Empirically, SpecRover solves 19.3% of issues on the full SWE-bench and 31.0% on SWE-bench Lite at an average cost of \$0.65 per issue. These results suggest that positioning an LLM as a high-level reviewer that reconciles specifications, tests, and natural language intent can substantially strengthen LLM-as-Judge, turning it into a calibrated confidence filter and explanation generator rather than a standalone oracle.
LLM-as-Judge components recast patch evaluation as a learned decision problem, letting repair systems trade broader exploration for higher precision at acceptance time. Across tree search, abstention and validation policies, and specification-guided review, they already improve solve rates, suppress noisy patches that merely satisfy weak tests, and provide textual justifications that help developers understand decisions. Remaining challenges include calibration, robustness to spurious cues, and alignment with stronger oracles, so future work should more closely couple judges with static analysis and regression testing, learn explicit abstention behavior, and treat these critics as conservative gatekeepers rather than standalone authorities.
@@ 7.3. Self-Controlled System
Self-controlled systems remove the handwritten skeleton altogether. One or more LLMs generate or revise the global plan at run time, spawn new sub-tasks when needed, and decide when to terminate, so the model, not the script, controls the entire outer loop. Representative examples are MAGIS (Tao et al., 2024), which distributes work to four specialized roles; SWE-Search (Antoniades et al., [n. d.]; Örwall, 2024), which couples LLMs with Monte Carlo tree search to debate and apply patches; and Learn-by-Interact (Hongjin et al., 2025), which lets a single model explore an environment, refine its strategy, and distill the trajectory into reusable instructions. With full control, these agents can tackle multi-file or cross-module defects but incur higher coordination costs and safety risks.
@@
SWE-Search combines Monte Carlo Tree Search with a multi-agent control loop in which GPT-4o iteratively searches the repository, plans edits, and debates candidate patches before applying a fix (Antoniades et al., [n. d.]; Örwall, 2024). On SWE-bench Lite, the system raises the solve rate to 31% with GPT-4o and delivers a 23% relative gain across five LLMs, while the Claude 3.5 Sonnet achieves 39% fixes at 2.7 issues per dollar. However, the search depth scales with inference-time computation and has been noted as “probably pretty expensive,” pointing to future work on cost-aware depth control and quicker repository traversal.
@@
Learn-by-Interact is a self-controlled agentic framework that lets an LLM explore a new software or web environment, plan its action sequence, execute those actions, and retrospectively distill the trajectory into reusable instructions, so the LLM, rather than a fixed script, decides every step of the outer loop (Hongjin et al., 2025). The authors fine-tune a Codestral-22B backbone on synthesized trajectories and also provide a training-free variant where Claude-3.5 Sonnet retrieves the distilled instructions at run time. On SWE-bench, Claude-3.5 Sonnet attains 60.0% pass@1 and improves pass@1 by up to 12.2pp over the baseline. The approach depends on accurate and up-to-date environment documentation; when documentation is sparse or changing, the data-synthesis phase can yield low-quality trajectories and incur substantial computational overhead.
@@
Self-controlled agents outperform simpler agents on public benchmarks and emerge with three lessons. First, explicit role separation or self-generated sub-goals narrows the search space without sacrificing autonomy. Second, stable feedback signals, such as full test suites or critical debates, prevent blind exploration. Third, careful cost management remains crucial because token budgets, tool latency, and coordination overhead can offset gains in accuracy. Continued progress on adaptive stopping rules, incremental context caching, and lightweight self-reflection is essential before these fully autonomous agents can meet the throughput and reliability demands of continuous integration.